# TransFR: Transferable Federated Recommendation with Adapter Tuning on Pre-trained Language Models

### v2

@@ Abstract
Federated recommendations (FRs), facilitating multiple local clients to collectively learn a global model without disclosing user private data, have emerged as a prevalent on-device service. In conventional FRs, a dominant paradigm is to utilize discrete identities to represent clients and items, which are then mapped to domain-specific embeddings to participate in model training. Despite considerable performance, we reveal three inherent limitations that can not be ignored in federated settings, i.e., non-transferability across domains, ineffectiveness in cold-start settings, and potential privacy violations during federated training. To this end, we propose a transferable federated recommendation model, TransFR, which delicately incorporates the general capabilities empowered by pre-trained models and the personalized abilities by fine-tuning local private data. Specifically, it first learns domain-agnostic representations of items by exploiting pre-trained models with public textual corpora. To tailor for FR tasks, we further introduce efficient federated adapter-tuning and test-time adaptation mechanisms, which facilitate personalized local adapters for each client by fitting their private data distributions. We theoretically prove the advantages of incorporating adapter tuning in FRs regarding both effectiveness and privacy. Through extensive experiments, we show that our TransFR surpasses several classic FRs on transferability.
@@ Introduction
Federated recommendations (FRs), owing to the benefits of efficient and secure distributed learning, have recently attracted a lot of interest both in academia and industry (Jiang et al. 2025). Empirically, it can be ingeniously generalized into a unified FR framework, which initially involves a user encoder and a fusion module in local client, and an item encoder in central server, as illustrated in Fig. 1. Many canonical works have expanded this research line in three directions (Li, He, and Song 2021). For instance, HPFL explores learning a powerful user encoder (Wu et al. 2021), RULE focuses on modeling efficient item embeddings (Chen et al. 2021), and Zhang et al. aims to design adaptive fusion modules (Zhang et al. 2023a). Subsequent research has built upon this foundation by adding textual data into FR fields to mitigate the data sparsity issue (Li et al. 2025c; Hu et al. 2025; Zhang et al. 2024a). For instance, IFedRec incorporates the textual descriptions of items based on user interactions to improve model performance (Zhang et al. 2024a). It is worth noting that all these federated models mainly utilize discrete identities (IDs) to represent items (textual data is only treated as side information), which have been an effective solution within single domains.
@@
Although ID-based FR models have been well established, they routinely suffer from the following multilevel challenges in cross-domain tasks. First, it lacks transferability across platforms, as item IDs are generally not shared in practical scenarios. To be specific, the same IDs across different domains may have different meanings, hence the embeddings trained with these IDs are domain-specific, which seriously limits the development of transferable FR models (Muhammad et al. 2020; Liu et al. 2021). Second, it fails to handle the cold-start issue due to the ID-based item embeddings. Specifically, a well-trained embedding highly relies on sufficient ID interactions, which cannot provide satisfactory results when users and items have few interactions (Zhang et al. 2024a). Third, it poses a privacy breach issue in FR settings since the item embeddings are trained directly with user private data (Zhang et al. 2024b). In summary, none of these FR models take into account all the issues, which are the primary challenges for real-world FRs.
@@
Given the challenges of existing work, we believe it is essential to develop a universal federated recommendation model that can effectively transfer to a new domain while ensuring unseen items can be handled well in cold-start settings. To achieve this, we can simply resort to the notion of retraining techniques to train the model from scratch with new domain-specific data (Zhang et al. 2020). Besides, we can augment side information or perturbations during the retraining process to alleviate the cold-start and privacy issues, respectively (Zhang et al. 2024a; Li, Song, and Fragouli 2020). However, such a straight solution brings a significant efficiency bottleneck since retraining the model involves multiple rounds of communication in federated learning (Guo et al. 2025; Li, He, and Song 2021; Zhang et al. 2023c), especially in resource-limited cross-device tasks. Moreover, retraining implies directly accessing to local raw data in the new domain, undoubtedly increasing the risk of user privacy leakage. Hence, it is imperative to design an efficient transferable FR framework, which can represent items in a domain-agnostic manner, following a privacy-preserving paradigm.
@@
Inspired by recent findings suggesting that ID features of items can be replaced by textual features (Li et al. 2025a), we present TransFR, a principled transferable FR framework built on pre-trained language models and federated adapter tuning techniques. Specifically, it represents items in a domain-agnostic way by purely utilizing fine-grained textual embeddings (as shown in the right panel of Fig. 1) instead of discrete item IDs. During this phase, language models like BERT are trained on publicly available text corpora on resource-rich servers. It can be understood that before the enactment of GDPR [^1] https://gdpr-info.eu/, these models could be trained with centrally collected data to achieve superior performance. After that, clients conduct efficient federated adapter tuning process, adapting it to the recommender task without transmitting raw data. This phase can be seen as post-GDPR period, adhering to legal restrictions on data stored at local clients. Finally, we introduce the test-time adaptation to further enhance the personalization of FR tasks. Overall, the main contributions of this work are listed as follows:
@@
  • We propose a unified FR model, TransFR, which can delicately exploits universal capabilities from pre-trained language models and the personalized abilities from fine-tuning techniques. To the best of our knowledge, this is the first work focusing on textual transferability in FRs.
@@
  • To achieve personalized FRs, we introduce a test-time adaptation module, with locally fitting a prediction head on private data distributions at inference time, effectively mitigating data heterogeneity across clients.
  • Extensive experiments on several datasets show the advantages of our model on cold-start effectiveness and transferability over state-of-the-art counterparts.
@@
In this section, we review two relevant areas to this work, i.e., federated recommendation (Yin et al. 2024) and transferable recommendation (Zang et al. 2022).
@@
Federated recommendation (FR) has emerged as a promising mobile service in recent years, due to its ability to preserve user privacy (Li et al. 2025b; Yi et al. 2025). FCF (Ammad-Ud-Din et al. 2019) and FedRec (Lin et al. 2020) are two pioneering works to utilize the discrete identities (IDs) to learn user/item embeddings. In general, there are three major components in the standard FR framework, i.e., user encoder, item encoder, and fusion module (Zhang et al. 2023a). As for the user encoder, Luo et al. proposed a personalized user encoder by combining the global and local clustered models (Luo, Xiao, and Song 2022). As for the item encoders, researchers have been working to address the privacy (Chen et al. 2025) and efficiency (Zhang et al. 2023b) issues associated with representing item embeddings with IDs. As for the fusion module, FedCIA proposed a similarity-based model fusion strategy to prevent information loss during the aggregation process (Han et al. 2025). Recent work, IFedRec (Zhang et al. 2024a), leveraged item attributes as side information to alleviate the cold-start issue in federated recommendation framework. Different from relying on coarse-grained IDs, our work represents items with fine-grained textual tokens to alleviate the transferability, cold-start and privacy issues of FRs.
@@
Transferable recommendation aims to improve target-domain performance by leveraging auxiliary data from other domains, mitigating data sparsity and cold-start issues (Zhang et al. 2025). Early approaches relied on explicit or implicit domain overlaps, using mapping functions for knowledge transfer (Man et al. 2017). For instance, EMCDR used a multi-layer perceptron for mapping to improve the robustness of target domain(Man et al. 2017). Distinct from the above work, subsequent efforts focus on learning universal representations with language models across different domains (Yuan et al. 2023; Xin et al. 2025). For example, UniSRec utilized the associated descriptions of items to learn transferable embeddings across different recommendation scenarios, showing the transferability in a cross-platform setting (Hou et al. 2022). Subsequent works aim to enhance transferability in federated settings. For example, FedCT enabled indirect cross-domain knowledge transfer but suffers from information loss due to spatial ID inconsistency (Liu et al. 2021). FedDCSR adopted disentangled ID-based embeddings for transfer, but may expose sensitive data during embedding transmission (Zhang et al. 2024b). FedCLR infered user interaction distributions in the source domain, reconstructs target-domain interactions via a generative model (Wang et al. 2025). Unlike prior methods that transfer item embeddings tied to private user data, our approach first learns transferable public parameters from open textual corpora, and then applies federated adapter tuning with private data, ensuring both privacy and transferability.
@@
Figure 2: The schematic diagram of our TransFR model, which includes hierarchical knowledge distillation, federated adapter-tuning and test-time adaptation mechanisms.
@@
We consider a standard federated recommendation scenario with $n$ users $\mathcal{U}=\{u\}$ and $m$ items =$\mathcal{I}=\{i\}$ on a central server. Each user $u$ holds local feedback data $\mathcal{D}_{u}$ consisting of interactions $(u,i,y_{ui})$, where $y_{ui}=1$ indicates user-item interaction. The goal is to predict $\hat{y}_{ui}$ for unseen items and recommend the top $K$ items to each user. In transferable recommendation, we define a source domain $S$ and multiple target domains $T=\{T_{1},\dots,T_{D}\}$, each with their own user/item sets. In the source domain (and similarly in target domains), there exists a user set $\mathcal{U}^{s}$ ($\mathcal{U}^{t}$) and an item set $\mathcal{I}^{s}$ ($\mathcal{I}^{t}$), comprising $|\mathcal{U}^{s}|$ ($|\mathcal{U}^{t}|$) users and $|\mathcal{I}^{s}|$ ($|\mathcal{I}^{t}|$) items. Our work focuses on cross-domain item recommendation for shared clients, utilizing domain-agnostic textual embeddings as an exploration in federated settings.
@@
Fig. 2 illustrates our framework. First, a central server pre-trains a large BERT model $\varphi_{\mathrm{BERT}}$ on public text data. To suit resource-limited devices, hierarchical knowledge distillation produces a smaller DistBERT $\varphi_{\mathrm{DistBERT}}$. DistBERT is then combined with an Adapter $\psi$ for efficient federated adapter-tuning, forming the Text Encoder $g=\psi\circ\varphi$. This encoder is deployed to user devices for local training. For personalization, a test-time adaptation module adds a local head $h_{u}$ per client, creating a personalized Item Encoder $\phi_{i}=h_{u}\circ g$. For the user representation $\mathbf{p}_{u}$ on local clients, we utilize local data to perform local training of the User Encoder $\phi_{u}$. Ultimately, we can utilize the fusion module $f(\cdot)$ to facilitate the prediction function for recommending item $i$ to user $u$, denoted as $\hat{y}_{ui}=f(\mathbf{p}_{u},\mathbf{q}_{i})$, where $\mathbf{q}_{i}$ denotes the produced embedding of item $i$. This process iterates until convergence.
@@
Hierarchical Knowledge Distillation. To enable efficient deployment on end-user devices, we introduce a hierarchical knowledge distillation (HKD) method tailored for BERT. HKD transfers knowledge from a larger teacher BERT model with $l$ Transformer layers to a smaller student model with $s$ layers ($l>s$), both sharing the same architecture except for depth. The goal is to reduce model size and speed up inference while maintaining accuracy by minimizing a defined distillation loss.
@@
where $h$ is the number of attention heads, and $\mathbf{A}_{i}\in\mathbb{R}^{L\times L}$ is the attention matrix of the $i$-th head, with $L$ as the input length. Inspired by studies showing BERT’s attention captures rich knowledge (Clark et al. 2019; Jiao et al. 2019), we use mean squared error loss to transfer knowledge from teacher to student. Besides attention, we also distill knowledge from the embedding layer, the objective is defined as:
@@
In TransFR, clients handle local optimization for federated adapter-tuning by updating user embeddings and uploading adapter gradients for global aggregation. They also perform test-time adaptation with local private data to personalize recommendations.
@@
Test-time Adaptation. Inspired by recent pioneering work, which indicates that the success of centralized machine learning is often attributable to the shared global feature representation, while the heterogeneity across clients or tasks predominantly manifests in the labels (LeCun, Bengio, and Hinton 2015). To mitigate this heterogeneity, we introduce the test-time adaptation (TTA) mechanism with a local prediction head $h_{u}(\cdot)$ on top of the Text Encoder $g(\cdot)$ for each client, which involves local training on the fine-tuned Item Encoder $\phi_{i}=h_{u}\circ g$ using private local data. This mechanism is designed to further fit the local data distributions, thereby facilitating personalized FRs. Specifically, we update the local prediction head using the local loss function $\mathcal{L}_{\mathrm{local}}$ in Eq.(6), detailed below:
@@
From a theoretical standpoint, this TTA strategy can be explained using domain adaptation theory (Ben-David et al. 2010). Let $P_{S}$ denote the global source distribution (on which the shared encoder $g$ is trained), and $P_{T}$ denote the client-specific target distribution. Then, the generalization error on the client domain can be bounded by:
@@
where $\epsilon_{T}$ and $\epsilon_{S}$ denote the expected prediction error on the target and source domains respectively, $\mathcal{D}_{\mathcal{H}}$ measures the divergence between the two distributions in hypothesis space $\mathcal{H}$, and $\lambda^{*}$ is the irreducible error of the optimal joint hypothesis. By updating $h_{u}$ locally to minimize $\epsilon_{T}$, the TTA mechanism reduces the distribution gap $\mathcal{D}_{\mathcal{H}}$ and improves the personalization capacity of the model.
@@
    incurs residual error increasing with the heterogeneity of $\{\mathbf{B}_{u}^{*}\}$. See Appendix for detailed derivation.
@@
[…]
@@
These results suggest that federated adapter tuning leaks significantly less information compared to full-model or embedding updates, due to its structural bottleneck and optional DP perturbation. For detailed proofs, see Appendix.
@@
RQ3: How do the specific components in our proposed method work for FR tasks, i.e., pre-training, federated adapter-tuning, and test-time adaptation?
RQ4: How do certain hyperparameters within our method impact the FR tasks, i.e., the number of layers in the adapter, the number of epochs in test-time adaptation module?
@@
To benchmark our model against baselines, we perform experiments with the Amazon (Ni, Li, and McAuley 2019) and Douban (Zhu et al. 2020) datasets. We selected three categories as distinct domains: Movies, Music, and Books. The statistics are summarized in Table 1 of the Appendix. To assess our model and baselines, we employ a leave-one-out evaluation approach and two prevalent evaluation metrics in ranking tasks: HR and NDCG.
@@
We adopt two kinds of benchmarks for overall comparisons, i.e., centralized models, i.e., CMF (Singh and Gordon 2008), EMCDR  (Man et al. 2017), DisenCDR (Cao et al. 2022), and recent federated approaches, i.e., FCF (Ammad-Ud-Din et al. 2019), IFedRec (Zhang et al. 2024a), FELLRec (Zhao et al. 2025), FedCT (Liu et al. 2021) and FedDCSR (Zhang et al. 2024b). More implementation details can be found in Appendix.
@@
In this section, we detail the experimental results on Amazon dataset.
@@
Table 3 aims to address RQ2, reporting HR@10 and NDCG@10 on Amazon dataset in cold-start settings, where models are evaluated on unseen items in the target domain. We derive several key observations. First, the federated baseline FCF struggles due to its ID-based item representation. IFedRec offers slight improvement by adding textual data but still relies on IDs, limiting its effectiveness. FedDCSR uses disentangled representations to transfer knowledge across domains, partially alleviating cold-start issues. Our method surpasses both ID-based and ID-mapping baselines by leveraging domain-agnostic textual embeddings and federated adapter tuning. Unlike coarse-grained ID embeddings, our textual representations capture item semantics more effectively and remain consistent across domains. This approach inherently addresses cold-start challenges.
@@
To address RQ3, we conduct an ablation study to assess the contribution of each component, i.e., pre-training (PT), federated adapter-tuning (FAT), and test-time adaptation (TTA), as shown in Fig.3 and Fig.4. We compare three variants, each removing one module: “w/o PT” replaces pre-training with random initialization, “w/o FAT” omits federated adapter tuning, and “w/o TTA” removes local head adaptation. As shown in Fig.3 and Fig.4, we observe the following: (1) Removing pre-training significantly degrades performance, aligning with prior findings (Chen et al. 2022); (2) Federated adapter-tuning yields moderate gains, validating the effectiveness of our Adapter module; (3) Test-time adaptation further boosts transferability, highlighting the role of the local head in personalized recommendation.
Table 4: Results of TransFR with differential privacy.
@@
We further incorporate the differential privacy to enhance privacy capability, with $\delta$ denoting noise intensity into adapters following PFedRec and FedRAP, which leads to only a slight performance decline on Amazon dataset, as shown in Table 4. This demonstrates that TransFR maintains competitive performance while keeping privacy protection.
Table 5: Ablation study comparison of pre-trained modules between DistBERT and Llmama-2 in terms of recommendation accuracy, memory overhead (MB), communication (commu.) overhead (MB), and inference time (s).
@@
Gain 2.70% 2.73% --20,982% --20,982% --400%
@@
Figure 6: Performance analysis of TransFR with different local epochs in test-time adaptation module.
@@
To address RQ4, we analyze the results with different hyper-parameters. As shown in Fig. 5 on Amazon dataset, increasing the number of hidden layers generally improves performance in the target domain, though some tasks (e.g., Music → Books, Books → Movies) show diminishing returns after a certain point. Thus, a moderate increase in hidden layers can enhance cross-device recommendation performance. We further examine the effect of local training epochs in test-time adaptation on Amazon dataset, as shown in Fig.6. Performance generally improves with more epochs, but overfitting may occur in scenarios with limited local data. Balancing effectiveness and efficiency, we set the number of local training epochs to 6 for all client devices.

### v3

@@ Abstract
Federated recommendations (FRs), facilitating multiple local clients to collectively learn a global model without disclosing user private data, have emerged as a prevalent on-device service. In conventional FRs, a dominant paradigm is to utilize discrete identities to represent clients and items, which are then mapped to domain-specific embeddings to participate in model training. Despite considerable performance, we reveal three inherent limitations that can not be ignored in federated settings, i.e., non-transferability across domains, ineffectiveness in cold-start settings, and potential privacy violations during federated training. To this end, we propose a transferable federated recommendation model, TransFR, which delicately incorporates the general capabilities empowered by pre-trained models and the personalized abilities by fine-tuning local private data. Specifically, it first learns domain-agnostic representations of items by exploiting pre-trained models with public textual corpora. To tailor for FR tasks, we further introduce efficient federated adapter-tuning and post-adaptation personalization, which facilitate personalized adapters for each client by fitting local private data. We theoretically prove the advantages of incorporating adapter tuning in FRs regarding both effectiveness and privacy. Through extensive experiments, we show that our TransFR surpasses state-of-the-art FRs on transferability.
@@ Introduction
Federated recommendations (FRs), owing to the benefits of efficient and secure distributed learning (Feng et al. 2018), have recently attracted a lot of interest both in academia and industry (Zhang et al. 2025; Du et al. 2024; Wu et al. 2022). Empirically, it can be ingeniously generalized into a unified FR framework (Zhang et al. 2024b, e; Chen et al. 2026), which initially involves a user encoder and a fusion module in local client, and an item encoder in central server, as illustrated in Fig. 1. Many canonical works have expanded this research line in three directions. For instance, HPFL explores learning a powerful user encoder (Wu et al. 2021), RULE focuses on modeling efficient item embeddings (Chen et al. 2021), and Zhang et al. aims to design adaptive fusion modules (Zhang et al. 2023a). Subsequent research has built upon this foundation by adding textual data into FR fields to mitigate the data sparsity issue (Li et al. 2025c; Hu et al. 2025). For instance, IFedRec incorporates the textual descriptions of items based on user interactions to improve model performance (Zhang et al. 2024a). It is worth noting that all these federated models mainly utilize discrete identities (IDs) to represent items, which have been an effective solution within single domains.
@@
Although ID-based FR models have been well established, they routinely suffer from the following multilevel challenges in cross-domain tasks. First, it lacks transferability across platforms, as item IDs are generally not shared in practical scenarios. To be specific, the same IDs across different domains may have different meanings, hence the embeddings trained with these IDs are domain-specific, which seriously limits the development of transferable FR models (Muhammad et al. 2020). Second, it fails to handle the cold-start issue due to the ID-based item embeddings. Specifically, a well-trained embedding highly relies on sufficient ID interactions, which cannot provide satisfactory results when users and items have few interactions (Zhang et al. 2024a). Third, it poses a privacy breach issue in FR settings since the item embeddings are trained directly with user private data (Zhang et al. 2024d). In summary, none of these FR models take into account all the issues, which are the primary challenges for real-world FRs.
@@
Given the challenges of existing work, we believe it is essential to develop a universal FR model that can effectively transfer to a new domain while ensuring unseen items can be handled well in cold-start settings. To achieve this, we can simply resort to the notion of retraining techniques to train the model from scratch with new domain-specific data (Zhang et al. 2020). Besides, we can augment side information or perturbations during the retraining process to alleviate the cold-start and privacy issues, respectively (Zhang et al. 2024a; Li, Song, and Fragouli 2020). However, such a straight solution brings a significant efficiency bottleneck since retraining the model involves multiple rounds of communication in federated learning (Guo et al. 2025; Feng et al. 2025), especially in resource-limited cross-device tasks. Moreover, retraining implies directly accessing to local raw data in the new domain, undoubtedly increasing the risk of user privacy leakage. Hence, it is imperative to design an efficient transferable FR framework, which can represent items in a domain-agnostic manner, following a privacy-preserving paradigm.
@@
Inspired by recent findings suggesting that ID features of items can be replaced by textual features (Li et al. 2025a), we present TransFR, a principled transferable FR framework built on foundation models and federated adapter tuning techniques. Specifically, it represents items in a domain-agnostic way by purely utilizing fine-grained textual embeddings (as shown in the right panel of Fig. 1) instead of discrete item IDs. During this phase, language models like BERT are trained on publicly available text corpora on resource-rich servers. It can be understood that before the enactment of GDPR [^1] https://gdpr-info.eu/, these models could be trained with centrally collected data to achieve superior performance. After that, clients conduct efficient federated adapter tuning process, adapting it to the recommender task without transmitting raw data. This phase can be seen as post-GDPR period, adhering to legal restrictions on data stored at local clients. Finally, we introduce post-adaptation personalization to further enhance the personalization of FR tasks. Overall, our main contributions are listed as follows:
@@
  • We propose a unified FR model, TransFR, which can delicately exploits universal capabilities from foundation models and the personalized abilities from fine-tuning techniques. To the best of our knowledge, this is the first work focusing on textual transferability in FRs.
@@
  • To achieve personalized FRs, we introduce a post-adaptation personalization module, with locally fitting a prediction head on private data distributions, effectively mitigating data heterogeneity across clients.
  • Extensive experiments on several datasets show the advantages of our model on effectiveness and transferability over state-of-the-art counterparts.
@@
In this section, we review two relevant areas to this work, i.e., federated recommendation (Zhang et al. 2025) and transferable recommendation (Zang et al. 2022).
@@
Federated recommendation (FR) has emerged as a promising mobile service in recent years, due to its ability to preserve user privacy (Li et al. 2025b; Yi et al. 2025; Chen et al. 2025). FCF (Ammad-Ud-Din et al. 2019) and FedRec (Lin et al. 2020) are two pioneering works to utilize the discrete identities (IDs) to learn user/item embeddings. In general, there are three major components in the standard FR framework, i.e., user encoder, item encoder, and fusion module (Zhang et al. 2023a). As for the user encoder, Luo et al. proposed a personalized user encoder by combining the global and local clustered models (Luo, Xiao, and Song 2022). As for the item encoders, researchers have been working to address the privacy (Li, Song, and Fragouli 2020) and efficiency (Zhang et al. 2023b) issues associated with representing item embeddings with IDs. As for the fusion module, FedCIA proposed a similarity-based model fusion strategy to prevent information loss during the aggregation process (Han et al. 2025). Recent work, IFedRec (Zhang et al. 2024a), leveraged item attributes as side information to alleviate the cold-start issue in FR framework. Different from relying on coarse-grained IDs, our work represents items with fine-grained textual tokens to alleviate the transferability, cold-start and privacy issues of FRs.
@@
Transferable recommendation aims to improve target-domain performance by leveraging auxiliary data from other domains, mitigating data sparsity issues (Zang et al. 2022; Bao et al. 2022; Zhang et al. 2024c). Early approaches relied on explicit or implicit domain overlaps, using mapping functions for knowledge transfer (Man et al. 2017). For instance, EMCDR used a multi-layer perceptron for mapping to improve the robustness of target domain(Man et al. 2017). Distinct from the above work, subsequent efforts focus on learning universal representations with language models across different domains (Yuan et al. 2023). For example, UniSRec utilized the associated descriptions of items to learn transferable embeddings across different scenarios, showing the transferability in a cross-platform setting (Hou et al. 2022). Subsequent works aim to enhance transferability in federated settings. For example, FedDCSR adopted disentangled ID-based embeddings for transfer, but may expose sensitive data during embedding transmission (Zhang et al. 2024d). FedCLR infered user interaction distributions in the source domain, reconstructs target-domain interactions via a generative model (Wang et al. 2025). Unlike prior methods that transfer item embeddings tied to private user data, our approach first learns transferable public parameters from open corpora, and then applies adapter tuning with private data, ensuring both privacy and transferability.
@@
Figure 2: The schematic diagram of TransFR, which includes hierarchical knowledge distillation, federated adapter-tuning and post-adaptation personalization (PAP).
@@
We consider a standard federated recommendation scenario with $n$ users $\mathcal{U}=\{u\}$ and $m$ items $\mathcal{I}=\{i\}$ on a central server. Each user $u$ holds local feedback data $\mathcal{D}_{u}$ consisting of interactions $(u,i,y_{ui})$, where $y_{ui}=1$ indicates user-item interaction. The goal is to predict $\hat{y}_{ui}$ for unseen items and recommend the top $K$ items to each user. In transferable recommendation, we define a source domain $S$ and multiple target domains $T=\{T_{1},\dots,T_{D}\}$, each with their own user/item sets. In the source domain (and similarly in target domains), there exists a user set $\mathcal{U}^{s}$ ($\mathcal{U}^{t}$) and an item set $\mathcal{I}^{s}$ ($\mathcal{I}^{t}$), comprising $|\mathcal{U}^{s}|$ ($|\mathcal{U}^{t}|$) users and $|\mathcal{I}^{s}|$ ($|\mathcal{I}^{t}|$) items. Our work focuses on cross-domain item recommendation for shared clients, utilizing domain-agnostic textual embeddings as an exploration in federated settings.
@@
Fig. 2 illustrates our framework. First, a central server pre-trains a large BERT model $\varphi_{\mathrm{BERT}}$ on public text data. To suit resource-limited devices, hierarchical knowledge distillation produces a smaller DistBERT $\varphi_{\mathrm{DistBERT}}$. DistBERT is then combined with an Adapter $\psi$ for efficient federated adapter-tuning, forming the Text Encoder $g=\psi\circ\varphi$. This encoder is deployed to user devices for local training. For personalization, a post-adaptation personalization module adds a local head $h_{u}$ per client, creating a personalized Item Encoder $\phi_{i}=h_{u}\circ g$. For the user representation $\mathbf{p}_{u}$ on local clients, we utilize local data to perform local training of the User Encoder $\phi_{u}$. Ultimately, we can utilize the fusion module $f(\cdot)$ to facilitate the prediction function for recommending item $i$ to user $u$, denoted as $\hat{y}_{ui}=f(\mathbf{p}_{u},\mathbf{q}_{i})$, where $\mathbf{q}_{i}$ denotes the produced embedding of item $i$. This process iterates until convergence.
@@
Hierarchical Knowledge Distillation. To enable efficient deployment on end-user devices, we introduce a hierarchical knowledge distillation (HKD) method tailored for BERT (Jiao et al. 2019). HKD transfers knowledge from a larger teacher BERT model with $l$ Transformer layers to a smaller student model with $s$ layers ($l>s$), both sharing the same architecture except for depth. The goal is to reduce model size and speed up inference while maintaining accuracy by minimizing a defined distillation loss.
@@
where $h$ is the number of attention heads, and $\mathbf{A}_{i}\in\mathbb{R}^{L\times L}$ is the attention matrix of the $i$-th head, with $L$ as the input length. Inspired by studies showing BERT’s attention captures rich knowledge (Clark et al. 2019), we use mean squared error loss to transfer knowledge from teacher to student. Besides attention, we also distill knowledge from the embedding layer, the objective is defined as:
@@
In TransFR, clients handle local optimization for federated adapter-tuning by updating user embeddings and uploading adapter gradients for global aggregation. They also perform post-adaptation personalization with local data to enhance personalization.
@@
Post-Adaptation Personalization. Inspired by recent pioneering work, which indicates that the success of centralized machine learning is often attributable to the shared global feature representation, while the heterogeneity across clients or tasks predominantly manifests in the labels (LeCun, Bengio, and Hinton 2015). To mitigate this heterogeneity, we introduce post-adaptation personalization (PAP) mechanism with a local prediction head $h_{u}(\cdot)$ on top of the Text Encoder $g(\cdot)$ for each client, which involves local training on the fine-tuned Item Encoder $\phi_{i}=h_{u}\circ g$ using private local data. This mechanism is designed to further fit the local data distributions, thereby facilitating personalized FRs. Specifically, we update the local prediction head using the local loss function $\mathcal{L}_{\mathrm{local}}$ in Eq.(6), detailed below:
@@
From a theoretical standpoint, this PAP strategy can be explained using domain adaptation theory (Ben-David et al. 2010). Let $P_{S}$ denote the global source distribution (on which the shared encoder $g$ is trained), and $P_{T}$ denote the client-specific target distribution. Then, the generalization error on the client domain can be bounded by:
@@
where $\epsilon_{T}$ and $\epsilon_{S}$ denote the expected prediction error on the target and source domains respectively, $\mathcal{D}_{\mathcal{H}}$ measures the divergence between the two distributions in hypothesis space $\mathcal{H}$, and $\lambda^{*}$ is the irreducible error of the optimal joint hypothesis. By updating $h_{u}$ locally to minimize $\epsilon_{T}$, the PAP mechanism reduces the distribution gap $\mathcal{D}_{\mathcal{H}}$ and improves the personalization capacity of the model.
@@
    incurs residual error increasing with the heterogeneity of $\{\mathbf{B}_{u}^{*}\}$.
@@
Let $\psi_{u}$ be the adapter parameters trained on user data $\mathcal{D}_{u}$, with bounded value range [C,C][-C,C] and quantization resolution $\delta$. The information leakage $I(\psi_{u};\mathcal{D}_{u})$ is upper bounded by:
@@
These results suggest that federated adapter tuning leaks significantly less information compared to full-model or embedding updates, due to its structural bottleneck and optional DP perturbation.
@@
RQ3: How do the specific components in our proposed method work for FR tasks, i.e., pre-training, federated adapter-tuning, and post-adaptation personalization?
RQ4: How do certain hyperparameters within our method impact the FR tasks, i.e., the number of layers in the adapter, the number of epochs in post-adaptation personalization?
@@
To benchmark our model against baselines, we perform experiments with the Amazon (Ni, Li, and McAuley 2019) dataset. We selected three categories as distinct domains: Movies, Music, and Books. Within each domain, we excluded interaction records lacking titles of items and applied filters to retain users with at least 20 interactions and items with a minimum of 30 interactions, in alignment with previous studies (He et al. 2017; Zhao et al. 2020). To assess our model and baselines, we employ a leave-one-out evaluation approach and two prevalent evaluation metrics in ranking tasks: HR and NDCG.
@@
We adopt two kinds of benchmarks for overall comparisons, i.e., centralized models, i.e., CMF (Singh and Gordon 2008), EMCDR  (Man et al. 2017), DisenCDR (Cao et al. 2022), and recent federated approaches, i.e., FCF (Ammad-Ud-Din et al. 2019), IFedRec (Zhang et al. 2024a), FELLRec (Zhao et al. 2025), FedCT (Liu et al. 2021) and FedDCSR (Zhang et al. 2024d).
@@
In this section, we detail the experimental results on Amazon dataset to verify the effectiveness of the proposed model.
@@
Table 3 aims to address RQ2, reporting HR@10 and NDCG@10 on Amazon dataset in cold-start settings, where models are evaluated on unseen items in the target domain. We derive several key observations. First, the federated baseline FCF struggles due to its ID-based item representation. IFedRec offers slight improvement by adding textual data but still relies on IDs, limiting its effectiveness. FedDCSR uses disentangled representations to transfer knowledge across domains, partially alleviating cold-start issues. Our method surpasses both ID-based and ID-mapping baselines by leveraging domain-agnostic textual embeddings and federated adapter tuning. Unlike coarse-grained ID embeddings, our textual representations capture item semantics more effectively and remain consistent across domains.
@@
To address RQ3, we conduct an ablation study to assess the contribution of each component, i.e., pre-training (PT), federated adapter-tuning (FAT), and post-adaptation personalization (PAP), as shown in Fig.3 and Fig.4. We compare three variants, each removing one module: “w/o PT” replaces pre-training with random initialization, “w/o FAT” omits federated adapter tuning, and “w/o PAP” removes local head adaptation. As shown in Fig.3 and Fig.4, we observe the following: (1) Removing pre-training significantly degrades performance, aligning with prior findings (Chen et al. 2022); (2) Federated adapter-tuning yields moderate gains, validating the effectiveness of our Adapter module; (3) Post-adaptation personalization further boosts transferability, highlighting the role of the local head in personalized recommendation.
@@
Table 4: Results of TransFR with differential privacy.
We further incorporate the differential privacy to enhance privacy capability, with $\delta$ denoting noise intensity into adapters, which leads to only a slight performance decline, as shown in Table 4. This shows that TransFR maintains competitive performance while keeping privacy protection.
@@
Gain 2.70% 2.73% -20,982% -20,982% -400%
Table 5: Ablation study comparison of pre-trained modules between DistBERT and Llama-2 in terms of recommendation accuracy, memory overhead (MB), communication (commu.) overhead (MB), and inference time (s).
Figure 6: Performance analysis of TransFR with different local epochs in post-adaptation personalization module.
@@
To address RQ4, we analyze results with different hyperparameters. As shown in Fig. 5, increasing the number of hidden layers generally improves performance in the target domain, though some tasks (e.g., Music → Books, Books → Movies) show diminishing returns after a certain point. Thus, a moderate increase in hidden layers can enhance cross-device recommendation performance. We further examine the effect of local training epochs in post-adaptation personalization on Amazon dataset, as shown in Fig.6. Performance generally improves with more epochs, but overfitting may occur in scenarios with limited local data. Balancing effectiveness and efficiency, we set the number of local training epochs to 6 for all client devices.